1. How would you prepare a liter of "carbonate buffer" at a pH of 10.07? You are provided with carbonic acid (H2CO3), sodium hydrogen carbonate (NaHCO3), and sodium carbonate (Na2CO3).

Please give an explanation.

2. Provide the mole ratio to support your answer. (Give your answer in the form base : acid.)
Answer #1

ph = 10.07

This is a buffer so:

pH = pKa2 + log(A-2/HA-)

Ka2 = 4.7 × 10−11

pKa = -log(Ka2) =-log(4.7 × 10−11)= 10.33

substitute

pH = pKa2 + log(A-2/HA-) = 10.33 +  log(A-2/HA-) = 10.07

Solve for ratio

log(A-2/HA-) = 10.07-10.33 = -0.26

CO3-2 / HCO3 -=0.5495

CO3-2 = 0.5495*HCO3

If you need V = 1L then

[Na2CO3] = 1M

[NaHCO3] = 0.5495 M

2)

mole ratio

bse: acid must be... 1 : 0.5495